Circular No. 7480

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dg IN MCG +1-1-29 C. Yu and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova (mag ∼ 17.8) on an unfiltered image taken on Aug. 22.5 UT, with confirmation coming from an earlier image taken on Aug. 20.5 (mag ∼ 18.5). SN 2000dg is located at α = 0^{h}6^{m}13^{s}.81, δ = +8^{o}53′17′′.5 (equinox 2000.0), which is 7′′.1 west and 0′′.8 south of the nucleus of MCG +1-1-29. A KAIT image of the same field taken on Aug. 11.4 showed nothing at the position of the new object (limiting mag ∼ 19.0).

SUNGRAZING COMETS Further to IAUC 7479, D. Hammer has provided measurements of 12

more Kreutz sungrazing comets detected by SOHO instruments (discovery positions given below with citations to MPECs containing the reduced measurements and orbits by B. G. Marsden). The comets were found in archival data by Hammer (C/1998 G8), M. Oates (the remaining 1997 Apr. and 1998 Apr. objects), and X. Leprette (C/1997 M4, C/1997 N3).

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 G3 Apr. 4.103  1^{h}00.^{m}5  + 4^{o}50′  2000-Q06
C/1997 G4  5.439  1 05.4  + 5 20  2000-Q06
C/1997 G5  5.826  1 06.4  + 5 28  2000-Q06
C/1997 G6  11.177  1 26.1  + 7 21  2000-Q06
C/1997 M4 June 17.685  5 40.8  +21 47  2000-Q06
C/1997 N3 July 4.670  6 49.2  +21 50  2000-Q06
Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 G5 Apr. 6.368  1^{h}07.^{m}9  + 5^{o}37′  2000-Q07
C/1998 G6  6.810  1 09.1  + 5 55  2000-Q07
C/1998 G7  7.491  1 11.0  + 5 54  2000-Q07
C/1998 G8  9.614  1 19.6  + 6 53  2000-Q08
C/1998 H3  26.496  2 20.0  +12 21  2000-Q07
C/1998 H4  27.941  2 25.4  +12 48  2000-Q07

SUPERNOVA 2000db IN NGC 3949 CCD R magnitudes measured by K. Hornoch, Lelekovice, Czech Rep.:

Aug. 11.830 UT, 13.7; 14.841, 13.6; 15.835, 13.7; 17.846, 13.7; 19.817, 13.7.

2000 August 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green